OpenStudy (anonymous):

How to rewrite the this as Ax+By=C f(x)=√3(x+√3)+2 can you help me solve this one with complete process on how you solve this thank you in advance

OpenStudy (larseighner):

If you mean: \( f(x)=\sqrt{3(x+√3)}+2\) it cannot be done, because it will necessarily involve the second power of f(x) which we are going to call y. But if you mean \(f(x)=√3(x+√3)+2\) then it is not too painful. First express f(x) as y \( \begin{align} y&=√3(x+√3)+2 \cr y- \sqrt{3}x &= \sqrt{3}\sqrt{3} + 2 \cr y - \sqrt{3} &= 5 \cr A&=1, \quad B= \sqrt{3}, \quad C=5\end{align}\)